OpenStudy (anonymous):

"Which of the following could be an example of a function with a domain [a, oo) and range [b, oo) where a>0 and b>0?" Will put the options in the comment. (And I'm not really asking for you to just give me the answer, I just really don't know how to do this.)

OpenStudy (anonymous):

A.\[f(x)=^{3}\sqrt{(x-b)}+a\]B.\[f(x)=\sqrt{x-a}+b\]C.\[f(x)=^{3}\sqrt{x+a}-b\]D.\[f(x)=\sqrt{x+a}-b\]